Circular No. 6231

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 S1

The following additional precise positions are available:

1995 UT  α_{2000}  δ_{2000}  Observer
Sept.17.80635 9^{h}03^{m}33.^{s}14 −2^{o}26′20′′.9 Tanaka
19.80703 9 10 44.81 −0 31 00.9 Sugie
19.80990 9 10 45.41 −0 30 50.4  _{”}
19.81406 9 10 46.40 −0 30 35.9  _{”}
20.74921 9 14 19.99 +0 25 31.5 Garradd
20.75298 9 14 20.85 +0 25 45.2  _{”}
20.75631 9 14 21.65 +0 25 57.5  _{”}
20.76024 9 14 22.54 +0 26 12.0  _{”}
20.76758 9 14 24.28 +0 26 39.1  _{”}
20.79120 9 14 29.78 +0 28 04.5  _{”}
20.79249 9 14 30.05 +0 28 10.2  _{”}
20.79383 9 14 30.29 +0 28 13.9  _{”}
20.80255 9 14 32.38 +0 28 38.4 Kobayashi
20.80329 9 14 32.49 +0 28 40.0 Kojima
20.80701 9 14 33.38 +0 28 54.8 Kobayashi
20.80728 9 14 33.39 +0 28 55.6 Kojima
20.81096 9 14 34.30 +0 29 09.1 Kobayashi
20.81139 9 14 34.34 +0 29 10.3 Kojima
20.81985 9 14 36.42 +0 29 41.6 Nakamura
20.82222 9 14 36.98 +0 29 50.3  _{”}
20.82465 9 14 37.55 +0 29 58.9  _{”}

M. Tanaka (Iwaki). 0.20-m f/1.5 Schmidt. Measurer M. Koishikawa. Com-

municated by A. Nakamura, Kuma, Ehime, Japan. A. Sugie (Dynic Astronomical Observatory). 0.60-m f/5.0 reflector + CCD.

Communicated by A. Nakamura. G. J. Garradd (Loomberah, N.S.W.). 0.25-m f/4.1 reflector + CCD. T. Kobayashi (Oizumi). 0.25-m f/4.4 reflector + CCD. T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6 reflector + CCD.

Kinked tail > 25′ long in p.a. 265^{o}. Communicated by A. Nakamura. A. Nakamura (Kuma Kogen). 0.60-m f/6.0 Ritchey-Chrétien + CCD.

Total visual magnitude estimate by K. Černis, Vilnius, Lithuania (0.12m refractor): Sept. 20.09 UT, 6.7.

1995 September 20 Daniel W. E. Green